OpenStudy (anonymous):

Help. Giving Out Medals ;) 4(x+3)<= 44 A. x<= 16 B. x>= 16 C. x<= 8 D. x >= 8

OpenStudy (anonymous):

Distribute 4 into the parenthesis. 4x + 12 <= 44 Subtract 12 to both sides, what's 44 - 12?

OpenStudy (anonymous):

Can you subtract that? @gxldmouth

OpenStudy (anonymous):

32 (:

OpenStudy (anonymous):

@iGreen. can i get some help on mine

OpenStudy (anonymous):

32 .

OpenStudy (misty1212):

i have an idea \[4(x+3)=44\] if you divide by 4 you can have \[x+3=11\] then subtract 3

OpenStudy (anonymous):

4(x+3)<= 44 4x + 12 < 44 44 - 12 = 32 32 / 4 = the answer to the question ! to find the > or < answer, simply add or subtract 1 to the final answer, and see if it is <= 44. do you understand ?

OpenStudy (anonymous):

Yep, that gives us: 4x < 32 Now divide 4 to both sides, what's 32 / 4? @gxldmouth

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Yep, so that gives you: x < 8 That's your answer..

OpenStudy (anonymous):

*<=